January 13, 2011 FDC meeting
- Production (David)
- Blue Crab status
- Electronics (Fernando, Chris, Roger)
- Engineering (Bill, David, Lubomir)
- Wire stringing FDC Log Book, page 549
- Movement control
- Full-scale prototype studies
- Tracking resolution FDC Log Book, page 546 (Simon)
Participants: Bill, Fernando, Eugene, Dave, Roger, Simon, and Lubomir.
- According to Tim, Dominion Power will have the power ready by next week. There were weather related delays. Dave: the electrical wiring inside the clean room is done, but data cables remain to be installed. Dave is is also busy with paperwork related to the production: quality plan, conduct of operation, clean room procedure, etc.
- Roger: answer from Allflex regarding our foil inspection report. The packing we suggested can be done. The most difficult is to avoid the stains. They will respond to all the questions in our report later this week.
- Roger expects e-mails from Fernando and Bill regarding the acceptance of the first article rigid-flex assemblies. Bill checked for matching to the cathode foil and frame, but he will check it also in the drawings. Chris will do some more measurements before the Fernando's approval. After the meeting Roger mentioned (based on Casey's information) that the contacts of the new rigid-flexes are not tinned, but they were tinned on the flexes for the full-scale prototype. We will correct this so that they are tinned in the mass production. Lubomir was concerned of the ground strip on the cathode foil touching the flex: due to the strong coupling it may attenuate the signals; this can be estimated.
- Fernando; the first article signal cables had a problem: one of the connectors had opposite polarity. He is contacting the company. Other than that the quality is very good.
- fADC125: according to Fernando there are some modifications that have to be done for the next fADC125 version by March 2011. Gerard will be able to work on this starting February.
- Wire stringing: this week we started wire stringing (first time at JLab!). After some trials it works well now: we have strung ~50 sense and ~5 field wires. Results for the position measurements on few wires are shown at FDC logbook page 549. For these measurements we used the laser sensor and the encoder. Results: the sigma of the deviation from the nominal value for the distance between the sense wires, is 6 microns (very good result!), while it is ~25 microns for the sense to filed wire distance. Full scan of the wire positions will be done after all they are installed. Bill: as mentioned before by Eugene we need to define a reference point (detectable by the laser sensor) on the frame to relate it to the wires, it can be put on the strongback. Bill is working also on the attachments for the camera and the magnet.
- The plan is to string all the wires, to solder them and do position and tension measurements (highest priority since we have never done this before). Then we will continue with the gluing of the PCB rings.
- Larry from Cross Computing (representing Parker) was at JLab end of last week working on low level software for the stepper motor controller. We are considering making a contract (for no more than one week work) to finish the job. The software must be open so that we can modify it.
Full-scale prototype tests
- Simon presented his results for the FDC resolutions with 40/60 Ar/CO2 gas mixture (page 584). Using Garfield smearing of the drift times in the MC simulations he demonstrates that the effective resolution is close the the one with Gaussian smearing with a sigma of 200 microns. Eugene had questions regarding details of the procedure and the Garfield simulations. Next step: generating smearing with Garfield with the magnetic field on.